OpenStudy (anonymous):

A coin is flipped, a die is rolled, and a card is drawn from a standard deck of cards. What is the probability of the following even. P(tails and 3 and king)?

OpenStudy (andriod09):

Coin=(1/2) die, i will assume it is 1-6 sides, = (3/6) deck of cards=(4/52)

OpenStudy (anonymous):

Do I add those together?

OpenStudy (andriod09):

idr sorry. :/ idt so though.

OpenStudy (anonymous):

Okay, thank you anyways !

OpenStudy (andriod09):

np!